What is the dimensional formula of electrical primitivity of free space × (electric field)^2

Grade:12

1 Answers

Deepak Kumar Shringi
5 years ago
Permittivity =[Ɛo] = [Charge]2 /([Force] [Distance]2) =[TA]2 / [MLT-2] [L]2 {[A] is the dimensional formula of electric charge} =[TA]2 / [ML3T-2] =[M-1L-3T(2+2)A2] = [M-1L-3T4A2]